# On Identification of Dynamic Treatment Regimes with Proxies of Hidden Confounders

### v1

@@ 6 Discussion
The current literature on estimating optimal dynamic treatment regimes in the presence of unmeasured confounding is sparse. We contribute to this literature by establishing identification assuming the existence of special proxy variables and solutions to certain integral equations. We provide closed-form solutions and empirical results in the categorical case. However, estimation and inference in the continuous case can be significantly more challenging, as it potentially involves solving ill-posed integral equations and is a promising avenue for future research. It would also be interesting to develop new highly efficient and robust semiparametric estimators that exhibit better worst-case performance than the simple plug-in estimators we examined in the simulation study. Finally, examining the implications of non-unique solutions to the integral equations (Zhang et al. (2023), Bennett et al. (2023)) could be of interest.
Acknowledgements
@@
Appendix A Proofs of Main Results
A.1 Known results
We first state some well-known results about identification in longitudinal observational studies under sequential randomization. First, recall that the joint potential outcome density is as follows:
@@ A.2 Proof of Theorem 1
The first and fourth equalities follow from basic probability algebra, the second and fifth from Assumption 1 ($A_{1}\perp\!\!\!\perp W_{1}\mid U_{0},Y_{0}$ and $A_{2}\perp\!\!\!\perp\overline{W}_{2}\mid\overline{U}_{1},\overline{Y}_{1},A_{1}$), the third and sixth from Assumption 2, and the final equality is by the g-formula, restated in the previous subsection in Equation 17. ∎
A.3 Proof of Lemma 1
@@
where the second equality is by Assumption 1, namely ${W}_{1}\perp\!\!\!\perp Z_{1}\mid A_{1},Y_{0},U_{0}$. Then we can write equation (6) as
@@
$=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)\sum_{u_{0}}f\left(w_{1}|y_{0},a_{1},u_{0}\right)f(u_{0}\mid y_{0},a_{1},z_{1})$
@@
$=\sum_{u_{0}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},u_{0}\right)f(u_{0}\mid y_{0},a_{1},z_{1})$
@@ A.5 Proof of Proposition 2
Appendix B Existence and uniqueness of bridge functions
B.1 Existence
@@
  4. (d)
    For any square integrable $g_{1}$, we have that
@@ B.2 Uniqueness
${E}[g_{1}(\overline{W}_{2})\mid\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}]=0\text{ implies }g_{1}(\overline{W}_{2})=0.$
@@
Appendix C General number of timepoints
@@
$E[g(\overline{U}_{k-1}\mid\overline{y}_{k-1},\overline{z}_{k},\overline{a}_{k}]\text{ implies }g(\overline{U}_{k-1})=0.$
@@
Under Assumptions 7 and 8, the following equation holds:
@@
If Assumption 7 and 9 hold, then $h_{KK},\ldots,h_{K1}$ solve the respective integral equations from Assumption 8 as well.
@@
Appendix D Over-identified Case
@@
Let $\infty>\min(|W_{t}|,|Z_{t}|)\geq|U_{t-1}|$ for $t=1,2$. Also, suppose Assumptions 1 and 10 hold. Then there exist solutions $h_{22}$, $h_{21}$, and $h_{11}$ solving the respective integral equations from Assumption 2 and Equation 5. In the special case where $|W_{t}|=|Z_{t}|=|U_{t-1}|$ for $t=1,2$, the conclusion of Proposition 2 holds.
@@
These results show that in the categorical confounding scenario, there exist $h_{22},h_{21},h_{11}$ that solve the equations from Assumption 2 and Equation 5. In the special case where the cardinality of the $W$, $U$, and $Z$ variables are the same, the Moore-Penrose pseudoinverses are simply inverses. Then by rearranging equations 12, 14, 16, the $h_{22}$, $h_{21}$, and $h_{11}$ vectors have unique representations
@@
Appendix E Additional details about the simulation
E.1 Oracle estimator
@@ E.2 NUCA estimator
E.3 Additional simulation results
We display simulation results for the proxy estimators with order 4,5,7, and 8 order interactions. Specifically, over the 1000 iterations, we plot histograms of the regrets of each of the methods alongside the regrets of the NUCA estimator. The results are depicted in Figure 4. For the most part, all seem to perform similarly, with the order 4 and 5 interaction estimators performing slightly worse than the order 7 and 8 interaction estimators.

### v2

@@ 6 Discussion
The current literature on estimating optimal dynamic treatment regimes in the presence of unmeasured confounding is sparse. Although we establish identification when proxy variables are measured and completeness conditions are met, there are many promising avenues for future research. Estimation and inference in the fully general case where cardinalities of variables are unrestricted can be significantly more challenging, as it potentially involves solving ill-posed integral equations. It would also be interesting to develop new highly efficient and robust semiparametric estimators that exhibit better worst-case performance than the simple plug-in estimators we examined in the simulation study. Finally, examining the implications of non-unique solutions to the integral equations (Zhang et al. (2023), Bennett et al. (2023)) could be of interest.
Acknowledgement
@@
Appendix A Section A: Proofs of main results
Known results
We first state some well-known results about identification in longitudinal observational studies under sequential randomization. First, recall that the joint potential outcome density can be written as follows:
@@ Appendix A Section A: Proofs of main results
The first and fourth equalities follow from basic probability algebra, the second and fifth from Assumption 1 ($A_{1}\perp\!\!\!\perp W_{1}\mid U_{0},Y_{0}$ and $A_{2}\perp\!\!\!\perp\overline{W}_{2}\mid\overline{U}_{1},\overline{Y}_{1},A_{1}$), the third and sixth from Assumption 2, and the final equality is by the g-formula, restated in the previous subsection. ∎
Proof of Lemma 1
@@
where the second equality is by Assumption 1, namely ${W}_{1}\perp\!\!\!\perp Z_{1}\mid A_{1},Y_{0},U_{0}$. Then we can write
@@
$=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)\sum_{u_{0}}f\left(w_{1}|y_{0},a_{1},u_{0}\right)f(u_{0}\mid y_{0},a_{1},z_{1}).$
@@
$=\sum_{u_{0}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},u_{0}\right)f(u_{0}\mid y_{0},a_{1},z_{1}).$
@@
Appendix B Section B: Existence and uniqueness of bridge functions
Existence
@@ Appendix B Section B: Existence and uniqueness of bridge functions
(iv) For any square integrable $g_{1}$, we have that
@@
${E}(g_{1}(\overline{W}_{2})\mid\overline{y}_{1},\overline{z}_{2},\overline{a}_{2})=0\text{ implies }g_{1}(\overline{W}_{2})=0.$
@@
Appendix C Section C: General number of timepoints
@@
$E\{g(\overline{U}_{k-1}\mid\overline{y}_{k-1},\overline{z}_{k},\overline{a}_{k}\}\text{ implies }g(\overline{U}_{k-1})=0.$
@@
Under Assumptions 9 and 10, the following equation holds:
@@
If Assumption 9 and 11 hold, then $h_{KK},\ldots,h_{K1}$ solve the respective integral equations from Assumption 10 as well.
@@
Appendix D Section D: Over-identified Case
@@
Let $\infty>\min(|W_{t}|,|Z_{t}|)\geq|U_{t-1}|$ for $t=1,2$. Also, suppose Assumptions 1 and 12 hold. Then there exist solutions $h_{22}$, $h_{21}$, and $h_{11}$ solving the respective integral equations from Assumption 2 and Equation 4. In the special case where $|W_{t}|=|Z_{t}|=|U_{t-1}|$ for $t=1,2$, the conclusion of Proposition 2 holds.
@@
These results show that in the categorical confounding scenario, there exist $h_{22},h_{21},h_{11}$ that solve the equations from Assumption 2 and Equation 4. In the special case where the cardinality of the $W$, $U$, and $Z$ variables are the same, the Moore-Penrose pseudoinverses are simply inverses. Then by rearranging equations 14, 16, 18, the $h_{22}$, $h_{21}$, and $h_{11}$ vectors have unique representations
@@
Section F: Treatment bridge case
Proof of Theorem 2
Proof.
By the discreteness of $Y_{1}$ and $Y_{2}$, we may rewrite $f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)=E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mid y_{0}]$. By the latent sequential randomization, we know that
$E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mid y_{0}]=E\Big{[}\frac{\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}}{P(A_{2}=a_{2}\mid A_{1},\overline{Y}_{1},\overline{U}_{1})P(A_{1}=a_{1}\mid Y_{0},{U}_{0})}\mid y_{0}\Big{]}.$
For a derivation of this fact, see Tsiatis et al. [2019] page 224. The proof now proceeds similarly to Theorem 2 of Ying et al. [2023].
$E[q_{12}\left(\overline{y}_{1},\overline{Z}_{2},\overline{a}_{2}\right)\mathbf{1}\{Y_{2}=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}\mid y_{0}]$
$=E[E[q_{12}\left(\overline{y}_{1},\overline{Z}_{2},\overline{a}_{2}\right)\mathbf{1}\{Y_{2}=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}\mid\overline{A}_{2},\overline{U}_{1},\overline{Y}_{1}]\mid y_{0}]$
$=E[\mathbf{1}\{Y_{2}=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}E[q_{12}\left(\overline{y}_{1},\overline{Z}_{2},\overline{a}_{2}\right)\mid\overline{A}_{2},\overline{U}_{1},\overline{Y}_{1}]\mid y_{0}]$
$=E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}\frac{E[q_{11}(Y_{0},Z_{1},A_{1})\mid A_{1},U_{0},Y_{0}]}{P(A_{2}=a_{2}\mid A_{1},\overline{U}_{1},\overline{Y}_{1})}\mid y_{0}]$
$=E[E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}\frac{E[q_{11}(Y_{0},Z_{1},A_{1})\mid A_{1},U_{0},Y_{0}]}{P(A_{2}=a_{2}\mid A_{1},\overline{U}_{1},\overline{Y}_{1})}\mid A_{1},\overline{U}_{1},\overline{Y}_{1}]\mid y_{0}]$
$=E[\mathbf{1}\{{A}_{1}={a}_{1},Y_{1}=y_{1}\}\frac{E[q_{11}(Y_{0},Z_{1},A_{1})\mid A_{1},U_{0},Y_{0}]}{P(A_{2}=a_{2}\mid A_{1},\overline{U}_{1},\overline{Y}_{1})}E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2}\}\mathbf{1}\{{A}_{2}={a}_{2}\}\mid A_{1},\overline{U}_{1},\overline{Y}_{1}]\mid y_{0}]$
$=E[\mathbf{1}\{{A}_{1}={a}_{1}\}E[q_{11}(Y_{0},Z_{1},A_{1})\mid A_{1},U_{0},Y_{0}]\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}=y_{1}\}\mid y_{0}]$
$=E[\mathbf{1}\{{A}_{1}={a}_{1}\}E[q_{11}(Y_{0},Z_{1},A_{1})\mid A_{1},U_{0},Y_{0}]\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mid y_{0}]$
$=E[\frac{1}{P(A_{1}=a_{1}\mid U_{0},Y_{0})}E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mathbf{1}\{{A}_{1}={a}_{1}\}\mid U_{0},Y_{0}]\mid y_{0}]$
$=E[\frac{1}{P(A_{1}=a_{1}\mid U_{0},Y_{0})}E[\mathbf{1}\{{A}_{1}={a}_{1}\}\mid U_{0},Y_{0}]E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mid U_{0},Y_{0}]\mid y_{0}]$
$=E[\mathbf{1}\{Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\}\mid y_{0}]$
The first equality is by iterated expectation. The second by independence of $\overline{Z}_{2}$ and $Y_{2}$ given $\overline{A}_{2},\overline{U}_{1},\overline{Y}_{1}$. The third is by consistency and definition of $q_{12}$. The fourth is by iterated expectation. The fifth pulls out constants. The sixth is by independence between $A_{2}$ and $Y_{2}(a_{1},a_{2})$ conditional on $A_{1},\overline{U}_{1},\overline{Y}_{1}$ and cancellation. The seventh is by consistency. The eighth is by definition of $q_{11}$ and iterated expectation. The ninth is by independence of $A_{1}$ from potential outcomes $(Y_{2}(a_{1},a_{2}),Y_{1}(a_{1}))$ given $U_{0},Y_{0}$. The last is from iterated expectation. ∎
Proof of Lemma 2
Proof.
The proof logic follows that of Theorem B.2 from the supplementary material of Ying et al. [2023] and is thus omitted. ∎
Appendix E Section E: Additional details about the simulation
Proxy estimator
The proxy estimator(s) for the categorical case in the simulation simply use plug-in estimates for all of the quantities according to Proposition 2. Explicitly,
$\widehat{d}_{2}^{proxy}\left(\overline{y}_{1},a_{1}\right)$
$\widehat{V}_{2}^{proxy}\left(a_{1},\overline{y}_{1}\right)$
Oracle estimator
@@ Appendix E Section E: Additional details about the simulation
Additional simulation results
For the methods oracle, NUCA, and proxy, we plot the histograms of regrets for each sample size in Figure 2. We plot the difference in regret between NUCA and proxy in Figure 3. In addition, we display simulation results for the proxy estimators with order 4,5,7, and 8 order interactions. Specifically, over the 1000 iterations, we plot histograms of the regrets of each of the methods alongside the regrets of the NUCA estimator. The results are depicted in Figure 4. For the most part, all seem to perform similarly, with the order 4 and 5 interaction estimators performing slightly worse than the order 7 and 8 interaction estimators.